OpenStudy (anonymous):

Let f(x) = (x2 + 3x – 4) and g(x) = (x + 4) a. Find f • g and state the domain. b. Find f/g and state the domain. please go through this step by step if you canand i dont really understand this

OpenStudy (anonymous):

domain for first one will be all real nos for second it is R-{-4}

OpenStudy (anonymous):

how did you get that though i dont understand it can you go threw it step by step

OpenStudy (anonymous):

for first it is direct since it is defined for all x second is bit hard ur f/g=((x2 + 3x – 4) / (x + 4)) now denominator will become zero in case of x=-4 which is not defined so we have to eliminate it from our domain set

OpenStudy (anonymous):

i dont see how you got -4 tho

OpenStudy (anonymous):

can u tell me what is in denominator

OpenStudy (anonymous):

i dont know what that is ... thats why i asked to explain it please

OpenStudy (anonymous):

x^2+3x-4*x+4=x^3+3x^2-4x+4x^2+12x-16=x^3+7x^2+8x-16., domain all real no. it z for a.